shift summary - day shift - 6/19/99

Ron Ray (rray@fnpxsr)
Sat, 19 Jun 1999 15:29:50 -0500

Shift Summary for DAY shift, 19/JUN/99

On Shift: TB, RR
JW, PS, JH, EB were around for parts of the shift to help with accesses.

Hours of beam delivered at all intensities: 1.0 of sporadic beam
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 10 min
Hours of special runs (laser scans, muon runs...): Laser scan - 45 min

Accounting of lost time:

Problems with CsI and ETOT system on our end. Ground fault which
prevented accelerator from running for most of the shift.

========================================================================
Runs ended this shift:
Run 13664 - Laser scan

Tapes brought to FCC this shift: KQ1001, ... None
(including tapes for online split)

========================================================================
Accesses during shift:

JW and PS make access to fix DPMT. JH and RR make access to investigate
ETOT problems.

========================================================================
Overview of shift:
E.g., calibrate CsI, E832 datataking

Continuation of nightmare from owl shift. A row of bad channels appear
on rib 13TE. This is an old problem which has resurfaced. A special
DPMT with an additional resistor is necessary to properly terminate
the L2+ trigger line on this rib. Julie and Peter Shanahan made an
access to modify the DPMT. They were successful and the problem now
appears to be gone.

While we were still in access, it was noticed that the ETOT system had
tripped. The buttons on the slow daq window were red, but no alarm was
generated. We tried to reset the system from upstairs, but one of the
power supplies did not come up. Jason was paged and he and Ron made
an access to look at the system. It appears that the temperature
interlock had tripped and shut down the system. One of the temperature
sensors appeared to have been moved, perhaps unintentionally during
work being done earlier. The probe was moved back to its correct
location and Jason tweaked the temperature thresholds a bit. All now
seems well.

The activated helium tubes were found to be empty. There was still pressure
on the output side of the regulator, so the helium bags were not affected.
Since no documentation exists for this cryptic system, it took some time
to finally open new tubes and ensure a continuous flow of helium.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

---------------------------------------------------------
Ron Ray - Fermilab - (630)840-8090 - rray@fnal.gov
---------------------------------------------------------